Let demand be given by Q = 150 - P + 2Y. This is the same for all problems of this type. Let r = 10%. Let Y = 50 across both periods. Let MC = 0. Let reserves = 400. Consider the basic two-period model. What is consumption of the resource in the future?

92.86

107.14

154.76

200

none of the above
Answer #1

Answer option 5) all wrong

Q= 150-P+2*50

Q= 250-P

P= 250-Q

Q1+Q2= 400

At eqm,

250-Q1 = (250-Q2)/(1+r)

(250-Q1)*1.1 = (250-Q2)

put Q1= 400-Q2

275-1.1*Q1 = 250-Q2

25= 1.1(400-Q2) - Q2

25= 440-2.1Q2

Q2*= 197.62

​​​​​​